OpenStudy (anonymous):

Gerry has two porcelain spheres. The radii of the spheres are in a ratio of 3 : 5. What is the ratio between the volumes of the two spheres? Answer 3 : 5 27 : 50 27 : 125 81 : 125

OpenStudy (apoorvk):

volume= (4/3)(\[\pi r^{3}\])  so the ratio of volumes of spheres of radii r' and r'' with ratio of radii r':r''=3:5 will be (r')^3:(r'')^3 so ans is 3^3:5^3 = 27:125